# Vortices and Factorization

### v1

@@ 1 Introduction, Stationary Vortex Configurations
$\frac{d\omega}{dt}=\frac{\partial\omega}{\partial t}+(v\nabla)\omega=0.$
@@
This system also has the 2d electrostatic interpretation: (5) is nothing, but an equilibrium condition of $N$ point electric charges ${\cal Q}_{1},\dots,{\cal Q}_{N}$ interacting pair-wise through 2d Coulomb (logarithmic) potential and placed in the homogeneous electric field $\bar{k}$. The equilibrium configuration is a critical point of the electrostatic energy
@@
where “c.c.” stands for the complex conjugate. Note that equilibrium configurations are special in the sense that system (5) does not have solution for generic parameters $k$ and ${\cal Q}_{i}$s. For instance, two-vortex ($N=2$) equilibrium exists only if $k\not=0$ and ${\cal Q}_{1}=-{\cal Q}_{2}$.
@@
Equilibrium patterns in non-uniform background flows will be considered in section 9. Sections 10 and 11 are devoted to periodic configurations (vortex streets).
@@ 2 Vortices, Polynomials and Rational Primitives
Details on Adler-Moser polynomials will be given in the following sections. Here, we note that these polynomials were first found by Burchnall and Chaundy [7], later re-discovered in context of the Hadamard problem [21] by Lagnese and Stellmacher, and then by Adler and Moser as polynomial $\tau$-functions corresponding to rational solutions of the KdV hierarchy of integrable PDEs [1]. Burchnall and Chaundy [7] built these polynomials as solutions of the Tkachenko equation (9) by considering it as a linear second order ODE for $p$, where coefficients of ODE are determined by some polynomial $P_{n}=q$. Then the requirement that its two linearly independent solutions $p=P_{n-1}$ and $p=P_{n+1}$ are both polynomial is equivalent to rationality in $z$ of two indefinite integrals (rational primitives)
@@
It turned out that, apart from the Adler-Moser case there is another case with $\Lambda=2$ when there exists an infinite chain of polynomial solutions (8) that can be built with help of rational primitives [9, 22].
@@ 3 $\Lambda=1$ and $\Lambda=2$ Chains: Main Sequences
In more detail: Since $p=z^{l}+\dots$ and $q=z^{m}+\dots$ from the highest symbol of (8) we get the diophantine equation connecting $l$ and $m$
@@
$\begin{array}[]{ll}q^{\prime}_{i+1}q_{i}-q_{i+1}q^{\prime}_{i}=(3i+1)p_{i},\\ p^{\prime}_{i}p_{i-1}-p_{i}p^{\prime}_{i-1}=\left(6i-1\right)q_{i}^{4},\end{array}$
@@
Summarizing, we emphasise again that, for $\Lambda=1$ and $\Lambda=2$, any no-common root solutions of the bi-linear equation (8) can be obtained by the recursive procedure described in this section. Therefore, the main sequences (23) or (25,24) determine all possible static patterns of a finite number of vortices with ${\cal Q}_{i}\in\{-1,1\}$ or ${\cal Q}_{i}\in\{-1,2\}$ respectively.
@@ 4 Darboux Transformations and Factorization Chains
In other words, we can obtain a new operator $\hat{H}$ and its eigenfunctions $\hat{\psi}(x;\lambda)$ from the old operator $H$ and its eigenfunctions $\psi(x;\lambda)$ corresponding to the same eigenvalues $\lambda$. New operator and eigenfunctions are obtained with the help of an eigenfunction $\varkappa=\psi(x;\eta)$ of the old operator $H$. The transformation $H\to\hat{H}$, $\psi\to\hat{\psi}$ given by (30,31) is called the Darboux Transformation.
@@
where $C_{i}$ are arbitrary constants. Using (35), we can built iteratively the chain (34). Note, that at each step of the chain a free parameter, which is a constant of integration, appears in (35).
@@
Here $\psi_{i}$ are $n$ eigenfunctions of the initial Schrodinger operator $H_{0}$ corresponding to the distinct eigenvalues $\eta_{1},\eta_{2},\dots,\eta_{n}$, i.e. $H_{0}\psi_{i}=\eta_{i}\psi_{i}$, $1\leq i\leq n$, called “seed functions”, and ${\cal W}$ stands for the Wronskian determinant. From (37), we get the Wronskian representation of the intertwining operator [^5] Simplest proof of the Crum theorem uses the fact that, according to (31,36), $T_{n}$ annihilates $n$ seed functions $\psi_{i}$, $i=1,\dots,n$, i.e. kernel of $T_{n}$ is spanned by these seed functions. Then, since $T_{n}=\partial_{z}^{n}+\dots$, we immediately come to (39).
@@ 5 Darboux Transformation and Equilibrium Configurations
As we mentioned above, sequences of the equilibrium configurations can be produced with help of Darboux transformations. To built them, we need to introduce the factorizable function corresponding to a static configuration (6)
@@ 6 Configurations Related to Bi-spectral Systems
Note that, for convenience, we set the normalization constants $C_{i}$ in (53) to values at which $P_{i}(z)$ are monic [^8] In the even case, this differs from the original normalization of [19], where same $C_{i}=2i+1$ were chosen for both families.. For “even” family, from (53,54) we get $P_{2}=z^{2}+t_{1}$. So, using (51), we obtain the sequence of length 2 depending on one parameter $t_{1}$:
@@ 7 KWCC Transformation and Terminating Configurations
Therefore, the chain (57) can be built by the iterative applications of following transformations
@@
Obviously, the KWCC transformations (59,60,61) generate all classes of configurations considered above, i.e. those related to the odd and even bi-spectral families as well as to the $\Lambda=2$ case.
@@
These are actually relations (21), but now $d_{i}$ are not necessarily degrees (15) or (17) of the main sequences. This is because (63, 64) are derived from the KWCC chain, rather than from the bi-linear equation (8). Because of this fact, any polynomial solution of (64) that starts from a static configuration (e.g. from a single vortex) also corresponds to a sequence of static configurations.
Apart from infinite main sequences of the $\Lambda=1$ and $\Lambda=2$ cases, recurrence relations (63,64) allow to find sequences of configurations of finite lengths, i.e. terminating configurations. The “even” bi-spectral family of section 6, i.e. solutions of (63,64) corresponding to $\gamma_{i}=1$, $d_{1}=1/2$, is a family of terminating configurations for $\Lambda=1$.
@@ 8 Translating Configurations and Baker-Akhieser Functions
A constructive way to write down $\psi_{j}$ is to use the Wronskian representation of $T_{j}$, given by (43, 44). Then, from (67) we get
@@ 9 Equilibria in Non-Homogenous Background Fields
where $R_{i}(z)$ are rational in $z$, we can apply several Darboux transformations resulting in various equilibrium configurations. According to the Darboux-Crum theorem (37,38), the results of compositions of $n$ transformations can be written in the Wronskian form
@@ 10 Equilibria of Vortex Streets - Odd Family
Similarly to vortex patterns on the plane that are related to the rational solutions of the KdV hierarchy or the odd bi-spectral family, the trigonometric soliton solutions of the hierarchy are related to the patterns on a cylinder [16, 23]. The latter can be shown by methods similar to those used in sections 5, 8.
@@
A family of the trigonometric-rational configurations can be generated by the chain of Darboux transformations (33) with $\eta_{i}=k_{i}^{2}$, where $0<k_{1}<k_{2}<\dots<k_{i}<\dots$ is a sequence of positive integers. The chain starts from the free Schrodinger operator $H_{0}=-\partial_{z}^{2}$ and it is built with the help of “seed functions” that are its periodic eigenfunctions
@@
Then, according to the Crum theorem (37,38), at $n$th step of the chain we have
@@
The class of configurations determined by (81,77) contains all possible translating vortex street patterns with ${\cal Q}_{i}=\pm 1$, as will be shown in section 12.
@@ 11 Equilibria of Vortex Streets - Even Family
For integers $a$ and $b$, (82) is a soliton solution of KdV at specific “times” (corresponding parameters $k_{i}$ and $\zeta_{i}$ in (78, 77) can be found e.g. in refs [6] or [15]) [^13] In general, sequence of $n$ Darboux transformations of the chain (33), where $A_{i}=\partial_{z}-(a+i)\cot z+(b+i)\tan z$, $i=0,1,\dots$., results in the shift of parameters $a\to a+n$, $b\to b+n$ in the potential (82).. In this case, $u_{0}$ and sequences of its Darboux transforms belong to the odd family considered in the previous section.
@@
From (83), with the help of the Wronskian formula (78), we can construct eigenfunctions of Darboux transforms $H_{n}=-\partial_{z}^{2}+u_{0}-2(\log\tau_{n})^{\prime\prime}$ of $H_{0}$. These eigenfunctions are factorizable, and therefore correspond to static vortex street configurations. Note that, for generic $a$ and $b$, only one of two linearly independent solutions of the Schrodinger equation $H_{0}\varkappa_{i}=\lambda_{i}\varkappa_{i}$ is factorizable. Therefore, in difference from the soliton-related configurations (78,77), configurations generated by (83) do not depend on free parameters others than $a$ and $b$. Similarly to systems considered in section 9, configurations with generic $a$ and $b$ are parametrized only by sequences of distinct integers $0<k_{1}<k_{2}<\dots<k_{n}$.
@@ 12 Time-Dependent Darboux Transformations, Calogero-Moser Systems and Locus Configurations
As we saw in section 3, all possible static configurations of finite number of vortices of strengths ${\cal Q}_{i}=\pm 1$ are given by roots of the Adler-Moser polynomials. In the present section we will demonstrate that sequences of soliton $\tau$-functions (77,78) define all possible vortex street patterns with ${\cal Q}_{i}=\pm 1$. To do this, we will introduce dynamical systems that are completely integrable in the $\Lambda=1$ case, and whose fixed points coincide with static and translating configurations of the point vortices. These systems are related to Darboux transforms of the time-dependent free Schrodinger (or heat) equation.
@@
$\frac{d^{2}z_{i}}{dt^{2}}=-\frac{(2a\pm 1)^{2}}{z_{i}^{3}}-\sum_{j:j\not=i}\left(\frac{8}{(z_{i}-z_{j})^{3}}+\frac{8}{(z_{i}+z_{j})^{3}}\right)$
@@
Initial equation $H_{0}\psi=\psi_{t}$ has quasi-polynomial solutions $\psi=z^{a}\prod_{i=1}^{l}(z^{2}-z_{i}(t)^{2})$, and the Darboux transforms built from these solutions are quasi-rational. When $a$ is half-integer, fixed points of system (94) are configurations of even bi-spectral family from section 6.
@@
The periodic analog of the Tkachenko equation (99) is of the same form as the Darboux chain equation (101). Then, taking the above fact and the fact that $p$ and $q$ are necessarily soliton $\tau$-functions (as follows from locus conditions) into account, we come to the conclusion that $p$ and $q$ are soliton $\tau$-functions related by Darboux transformation. Therefore, all possible static street patterns consisting of vortices of two species of opposite strengths are given by (77,78).
For translating streets we use arguments similar to those of the rational case and come to the conclusion that the complete set of translating street patterns with ${\cal Q}_{i}=\pm 1$ is determined by (77,81). It is worth mentioning that from the soliton theory it follows that here the relationship between $p$ and $q$ holds
$q=p(z,\sigma_{1},\dots,\sigma_{n}),\quad p=p(z,\zeta_{1},\dots,\zeta_{n}),\quad(k+{\rm i}k_{j})e^{2{\rm i}\sigma_{j}}=(k-{\rm i}k_{j})e^{2{\rm i}\zeta_{j}},$
where $p=\tau_{n}$ is given by (77,78). Without going into much details, we note that the above relationship can be derived using the Sato formula for the Baker-Akhieser function of the KP hierarchy and the correspondence between Hirota and Wronskian representations of the $n$-soliton $\tau$-functions.

### v2

@@ 1 Introduction, Stationary Vortex Configurations
$\frac{d\omega}{dt}=\frac{\partial\omega}{\partial t}+(\mathbf{v}\nabla)\omega=0.$
@@
This system also has the 2d electrostatic interpretation: (5) is nothing, but an equilibrium condition of $N$ point electric charges ${\cal Q}_{1},\dots,{\cal Q}_{N}$ interacting pair-wise through 2d Coulomb (logarithmic) potential and placed in the homogeneous electric field $\bar{k}$. The equilibrium configuration is a stationary point of the electrostatic energy
@@
where “c.c.” stands for the complex conjugate.
@@
This system of equations is invariant under rigid motions and scaling of the plane. From invariance of stationary value of electrostatic energy under scaling of the plane it follows that the system has solutions only if $\sum_{i<j}Q_{i}Q_{j}=0$. O’Neil showed [31] that for almost all values of $Q_{i}$, satisfying the above condition, there are exactly $(N-2)!$ distinct configurations, i.e. modulo rigid motions and scaling of the plane, distinct solutions of (6).
In the case $k\not=0$ necessary and sufficient condition of existence of solutions is $\sum_{i}{\cal Q}_{i}=0$. For almost all values of ${\cal Q}_{i}$ satisfying this condition number of distinct, modulo translations of the plane, solutions of (5) is $(N-1)!$.
There are, however, “resonant” values of ${\cal Q}_{i}$, where number of configurations is infinite. This paper mainly deals with such resonant cases. Sections 10 and 11 are devoted to periodic (“vortex street”) versions of such systems.
We also briefly consider non-resonant cases in relationship with factorization method. These are equilibrium patterns in non-uniform background flows from section 9.
@@ 2 Vortices, Polynomials and Rational Primitives
Details on Adler-Moser polynomials will be given in the following sections. Here, we note that these polynomials were first found by Burchnall and Chaundy [7], later re-discovered in context of the Hadamard problem [23] by Lagnese and Stellmacher, and then by Adler and Moser as polynomial $\tau$-functions corresponding to rational solutions of the KdV hierarchy of integrable PDEs [1]. Burchnall and Chaundy [7] constructed these polynomials as solutions of the Tkachenko equation (9) by considering it as a linear second order ODE for $p$, where coefficients of ODE are determined by some polynomial $P_{n}=q$. Then the requirement that its two linearly independent solutions $p=P_{n-1}$ and $p=P_{n+1}$ are both polynomial is equivalent to rationality in $z$ of two indefinite integrals (rational primitives)
@@
It turned out that, apart from the Adler-Moser case there is another case with $\Lambda=2$ when there exists an infinite chain of polynomial solutions (8) that can be constructed with help of rational primitives [9, 24].
@@ 3 $\Lambda=1$ and $\Lambda=2$ Chains: Main Sequences
In more detail: Since $p=z^{l}+\dots$ and $q=z^{m}+\dots$ from the leading term of (8) we get the diophantine equation connecting $l$ and $m$
@@
$\begin{array}[]{ll}q^{\prime}_{i+1}q_{i}-q_{i+1}q^{\prime}_{i}=(3i+1)p_{i},\\ p^{\prime}_{i}p_{i-1}-p_{i}p^{\prime}_{i-1}=\left(6i-1\right)q_{i}^{4},\end{array}$ (24)
@@
Summarizing, we emphasise again that, for $\Lambda=1$ and $\Lambda=2$, any no-common root solutions of the bi-linear equation (8) can be obtained by the recursive procedure described in this section. Therefore, the main sequences (23) or (26,25) determine all possible static patterns of a finite number of vortices with ${\cal Q}_{i}\in\{-1,1\}$ or ${\cal Q}_{i}\in\{-1,2\}$ respectively.
@@ 4 Darboux Transformations and Factorization Chains
In other words, we can obtain a new operator $\hat{H}$ and its eigenfunctions $\hat{\psi}(z;\lambda)$ from the old operator $H$ and its eigenfunctions $\psi(z;\lambda)$ corresponding to the same eigenvalues $\lambda$. New operator and eigenfunctions are obtained with the help of an eigenfunction $\varkappa=\psi(z;\eta)$ of the old operator $H$. The transformation $H\to\hat{H}$, $\psi\to\hat{\psi}$ given by (31,32) is called the Darboux Transformation.
@@
where $C_{i}$ are arbitrary constants. Using (36), we can construct iteratively the chain (35). Note, that at each step of the chain a free parameter, which is a constant of integration, appears in (36).
@@
Here $\psi_{i}$ are $n$ eigenfunctions of the initial Schrodinger operator $H_{0}$ corresponding to the distinct eigenvalues $\eta_{1},\eta_{2},\dots,\eta_{n}$, i.e. $H_{0}\psi_{i}=\eta_{i}\psi_{i}$, $1\leq i\leq n$, called “seed functions”, and ${\cal W}$ stands for the Wronskian determinant. From (38), we get the Wronskian representation of the intertwining operator [^5] Simplest proof of the Crum theorem uses the fact that, according to (32,37), $T_{n}$ annihilates $n$ seed functions $\psi_{i}$, $i=1,\dots,n$, i.e. kernel of $T_{n}$ is spanned by these seed functions. Then, since $T_{n}=\partial_{z}^{n}+\dots$, we immediately come to (40).
@@ 5 Darboux Transformation and Equilibrium Configurations
As we mentioned above, sequences of the equilibrium configurations can be produced with help of Darboux transformations. To construct them, we need to introduce the factorizable function corresponding to a static configuration (6)
@@ 6 Configurations Related to Bi-spectral Systems
Note that, for convenience, we set the normalization constants $C_{i}$ in (54) to values at which $P_{i}(z)$ are monic [^8] In the even case, this differs from the original normalization of [17], where same $C_{i}=2i+1$ were chosen for both families.. For “even” family, from (54,55) we get $P_{2}=z^{2}+t_{1}$. So, using (52), we obtain the sequence of length 2 depending on one parameter $t_{1}$:
@@ 7 KWCC Transformation and Terminating Configurations
Therefore, the chain (58) can be constructed by the iterative applications of following transformations
@@
Obviously, the KWCC transformations (60,61,62) generate all classes of configurations considered above, i.e. those related to the odd and even bi-spectral families as well as to the $\Lambda=2$ case.
@@
These are actually relations (21), but now $d_{i}$ are not necessarily degrees (15) or (17) of the main sequences. This is because (64, 65) are derived from the KWCC chain, rather than from the bi-linear equation (8). Because of this fact, any polynomial solution of (65) that starts from a static configuration (e.g. from a single vortex) also corresponds to a sequence of static configurations.
Apart from infinite main sequences of the $\Lambda=1$ and $\Lambda=2$ cases, recurrence relations (64,65) allow to find sequences of configurations of finite lengths, i.e. terminating configurations. The “even” bi-spectral family of section 6, i.e. solutions of (64,65) corresponding to $\gamma_{i}=1$, $d_{1}=1/2$, is a family of terminating configurations for $\Lambda=1$.
@@ 8 Translating Configurations and Baker-Akhieser Functions
A constructive way to write down $\psi_{j}$ is to use the Wronskian representation of $T_{j}$, given by (44, 45). Then, from (68) we get
@@ 9 Equilibria in Non-Homogenous Background Fields
where $R_{i}(z)$ are rational in $z$, we can apply several Darboux transformations resulting in various equilibrium configurations. According to the Darboux-Crum theorem (38,39), the results of compositions of $n$ transformations can be written in the Wronskian form
@@ 10 Equilibria of Vortex Streets - Odd Family
Similarly to vortex patterns on the plane that are related to the rational solutions of the KdV hierarchy or the odd bi-spectral family, the trigonometric soliton solutions of the hierarchy are related to the patterns on a cylinder [19, 21, 25]. The latter can be shown by methods similar to those used in sections 5, 8.
@@
A family of the trigonometric-rational configurations can be generated by the chain of Darboux transformations (34) with $\eta_{i}=k_{i}^{2}$, where $0<k_{1}<k_{2}<\dots<k_{i}<\dots$ is a sequence of positive integers. The chain starts from the free Schrodinger operator $H_{0}=-\partial_{z}^{2}$ and it is constructed with the help of “seed functions” that are its periodic eigenfunctions
@@
Then, according to the Crum theorem (38,39), at $n$th step of the chain we have
@@
The class of configurations determined by (82,78) contains all possible translating vortex street patterns with ${\cal Q}_{i}=\pm 1$, as will be shown in section 12.
@@ 11 Equilibria of Vortex Streets - Even Family
For integers $a$ and $b$, (83) is a soliton solution of KdV at specific “times” (corresponding parameters $k_{i}$ and $\zeta_{i}$ in (79, 78) can be found e.g. in refs [6] or [18]) [^13] In general, sequence of $n$ Darboux transformations of the chain (34), where $A_{i}=\partial_{z}-(a+i)\cot z+(b+i)\tan z$, $i=0,1,\dots$., results in the shift of parameters $a\to a+n$, $b\to b+n$ in the potential (83).. In this case, $u_{0}$ and sequences of its Darboux transforms belong to the odd family considered in the previous section.
@@
From (84), with the help of the Wronskian formula (79), we can construct eigenfunctions of Darboux transforms $H_{n}=-\partial_{z}^{2}+u_{0}-2(\log\tau_{n})^{\prime\prime}$ of $H_{0}$. These eigenfunctions are factorizable, and therefore correspond to static vortex street configurations. Note that, for generic $a$ and $b$, only one of two linearly independent solutions of the Schrodinger equation $H_{0}\varkappa_{i}=\lambda_{i}\varkappa_{i}$ is factorizable. Therefore, in difference from the soliton-related configurations (79,78), configurations generated by (84) do not depend on free parameters others than $a$ and $b$. Similarly to systems considered in section 9, configurations with generic $a$ and $b$ are parametrized only by sequences of distinct integers $0<k_{1}<k_{2}<\dots<k_{n}$.
@@ 12 Time-Dependent Darboux Transformations, Calogero-Moser Systems and Locus Configurations
As we saw in section 3, all possible static configurations of finite number of vortices of strengths ${\cal Q}_{i}=\pm 1$ are given by roots of the Adler-Moser polynomials. In the present section we will demonstrate that sequences of soliton $\tau$-functions (78,79) define all possible vortex street patterns with ${\cal Q}_{i}=\pm 1$. To do this, we will introduce dynamical systems that are completely integrable in the $\Lambda=1$ case, and whose fixed points coincide with static and translating configurations of the point vortices. These systems are related to Darboux transforms of the time-dependent free Schrodinger (or heat) equation.
@@
$\frac{d^{2}z_{i}}{dt^{2}}=\frac{1-(2a\pm 1)^{2}}{z_{i}^{3}}-\sum_{j:j\not=i}\left(\frac{8}{(z_{i}-z_{j})^{3}}+\frac{8}{(z_{i}+z_{j})^{3}}\right)$
@@
Initial equation $H_{0}\psi=\psi_{t}$ has quasi-polynomial solutions $\psi=z^{a}\prod_{i=1}^{l}(z^{2}-z_{i}(t)^{2})$, and the Darboux transforms constructed from these solutions are quasi-rational. When $a$ is half-integer, fixed points of system (95) are configurations of even bi-spectral family from section 6.
@@
The periodic analog of the Tkachenko equation (100) is of the same form as the Darboux chain equation (102). Then, taking the above fact and the fact that $p$ and $q$ are necessarily soliton $\tau$-functions (as follows from locus conditions) into account, we come to the conclusion that $p$ and $q$ are soliton $\tau$-functions related by Darboux transformation. Therefore, all possible static street patterns consisting of vortices of two species of opposite strengths are given by (78,79).
For translating streets we use arguments similar to those of the rational case and come to the conclusion that the complete set of translating street patterns with ${\cal Q}_{i}=\pm 1$ is determined by (78,82). It is worth mentioning that from the soliton theory it follows that here the relationship between $p$ and $q$ holds
$q=\tau_{n}(z,\sigma_{1},\dots,\sigma_{n}),\quad p=\tau_{n}(z,\zeta_{1},\dots,\zeta_{n}),\quad(k+{\rm i}k_{j})e^{2{\rm i}\sigma_{j}}=(k-{\rm i}k_{j})e^{2{\rm i}\zeta_{j}},$
where $\tau_{n}$ is given by (78,79). Without going into much details, we note that the above relationship can be derived using the Sato formula for the Baker-Akhieser function of the KP hierarchy and the correspondence between Hirota and Wronskian representations of the $n$-soliton $\tau$-functions (for introduction to theory of solitons see e.g. [39]).